How many odd numbers of at most three digits can be formed using the digits 0,1,2,3,4,5 without repetitions?
It says the answer is 63. I have no clue how to get this answer.. Could I get some help please?

Answer by Alan3354(69443)

How many odd numbers of at most three digits can be formed using the digits 0,1,2,3,4,5 without repetitions?
It says the answer is 63.
===================
The units digit has to be 1, 3 or 5 to be odd, that's 1 of 3
------
Assuming leading zeroes are not allowed, eg, 025, (tho it should be specified, we should not have to make assumptions), the 100's digit for 3 digit numbers must be 1,2,3,4,or 5, but one of the odd numbers is used, so it's 1 of 4 for the 100's digit.
==================
That's 3*4 = 12 possible 3 digit odd numbers.
------------
There are 3 possible 1 digit numbers, 1, 3 & 5
-----------
For 2 digit numbers:
The units is 1 of 3.
The tens digit is 1 of 6
3*6 = 48
****************
48 + 12 + 3 = 63